OpenStudy (anonymous):

Your friend runs up to you, scared that he is not ready for the upcoming quadratics test. To help him study, you will create four different quadratic functions. Then demonstrate to him how to rewrite each function as a group of factors, if possible. The function f(x) is a difference of squares. The function g(x) is a sum of squares. The function h(x) is a perfect square trinomial. The function j(x) can only have a GCF factored out of it.

OpenStudy (anonymous):

Difference of squares: (a+b)(a-b) = a^2 - b^2 Sum of Squares: (a+b)(a+b) = a^2 + 2ab + b^2 Perfect Square Trinomial: x^2 + 4x - 5 = (x+5)(x-1) GCF: 16x^2 + 8x + 32 = 8(2x^2 + x + 4)

OpenStudy (anonymous):

How would I factor as a group the equations you gave me